# Unsupervised Representation Learning from Sparse Transformation Analysis

### v2

@@ 1 Introduction
One early line of work in representation learning focused on ideas of redundancy reduction, believing that biological neural systems would naturally strive for an efficient code due to competitive pressures [barlow1961possible]. Building on this idea, the principles of sparsity and statistical independence of coding dimensions emerged as guidelines for learning such maximally efficient codes, eventually resulting in the frameworks of sparse coding [olshausen1997sparse] and independent component analysis [comon1994independent]. Inspired by the fact that natural intelligence is embedded in a world where physical laws restrict observations to sequences of smooth transformations, these ideas of efficiency and sparsity were extended to include temporal dimensions. A seminal example is Slow Feature Analysis [wiskott2002slow], a learning framework which assumes that individual latent variables are likely to change slowly over time. Models adhering to these principles were shown to learn invariances directly from data and uncover underlying generative factors if those factors had similar slow dynamics. Recent work has further shown that natural videos follow a specific sparse transition structure, meaning that the set of generative factors which describe a given input sequence is mostly constant over time with sparse transitions between which factors are active. Klindt et al. [klindt2021towards] then demonstrated that by building a model which incorporates this structure into its prior, it is possible to provably learn the true generative factors of video data in an unsupervised manner. While differing in implementation and methodology, all these frameworks appear in some sense to share the goals of learning meaningful, interpretable, ‘disentangled’, and controllable latent codes such that specific directions in the latent space corresponded to the independent factors which were responsible for generating the input data distribution.
@@
More recently, the concept of equivariance has emerged as a mathematical framework for learning highly structured and thereby ‘controllable’ latent representations [cohen2016group]. Specifically, equivariant neural networks are built to explicitly respect the symmetries of the input domain in their output space. In such models, there are known predictable output transformations for given input transformations of interest. To date, these concepts form the foundation of some of the most precisely formalized definitions of ‘disentanglement’ in the literature [higgins2018towards, cohen2015transformation]. In prior work however, there is a relatively sharp divide between equivariant neural networks and models which are focused on disentanglement. Specifically, it is currently only known how to build networks which are equivariant with respect to known transformations which have a mathematical group structure. This includes traditional coordinate symmetries [cohen2017steerable, hoogeboom2022equivariant], but is severely limiting when considering the types of natural image transformations that are typically explored in disentangled representation learning. One line of research has aimed to bridge this divide by learning ‘approximately equivariant’ models which are intended to learn these types of structured representations directly from data itself [keller2021topographic, song2023flow]. However the vast majority of the models in this domain require at least some form of weak supervision of segmented sequences with only single transformations being observed.
@@
In this paper, we introduce a new modeling framework, denoted Sparse Transformation Analysis (STA), which takes inspiration from these foundational representation learning approaches, thereby yielding what we argue to be a uniquely structured yet flexible latent space which aligns with natural data statistics. Fig. 1 depicts the overview of our STA. The framework requires no supervision of input sequences, assuming only that the observed transformations from one timestep to the next match a sparse transition structure similar to that observed by [klindt2021towards]. Specifically, STA takes a generative modeling approach, asserting that generative factors should be represented by distributions over latent variables, and that these distributions should flow smoothly in the latent space in concert with the smooth flow of observations in the world. Furthermore, the framework posits that this flow should not be arbitrary, but can be represented as a sparse combination of learned flow field primitives. In alignment with notions of disentanglement and approximate equivariance, these flow field primitives can be seen as directions in latent space which correspond to observed input transformations. Unlike [klindt2021towards], STA allows for highly flexible latent dynamics for each transformation ‘direction’ by parameterizing each transformation’s flow field through the Helmholtz decomposition as a combination of curl-free and divergence-free components. The specific sparse combination of flow fields which are used to transform the latent distribution from one time-step to the next is treated as an unobserved latent variable with a multivariate history-dependent spike and slab prior [mitchell1988bayesian], and inferred simultaneously with the other latent variables through amortized variational inference. In this way, this approach can be seen to combine ideas of sparse coding, slow feature analysis, and approximate equivariance, while still allowing deep neural network feature extractors to be leveraged in a relatively unconstrained manner.
@@
In the following, we will demonstrate that this framework yields the state of the art in unsupervised approximate equivariance, as quantified through a measured equivariance error, and further that our method yields the highest likelihood on the test set in the unsupervised setting. As desired, we additionally observe that the model automatically learns to separate observed transformations into independent flow fields, and that these latent flows can be flexibly combined or switched during traversal. By tuning the magnitude of these flow fields, our model also has precise control of the transformation speed. Further, when slightly modifying the spike component to two separate controls, our method can learn to segregate latent symmetries and invariances into the two distinct components of the vector fields. The decomposition of latent flow fields is coherent with the categories of input transformations. Besides simple experiments on toy datasets, we further validate our STA on use cases of real-world video analysis, including movements of robot arms [nie2020semi], lighting changes of indoor scenes [nie2020semi], behavior videos of social agents [sun2021multi], and ego-centric autonomous driving videos [cordts2016cityscapes]. Our method can identify a wide range of independent motions in the video sequences. Ultimately, we present this model as a natural next step in the development of unsupervised approximately equivariant representation learning algorithms. Code is publicly available at https://github.com/KingJamesSong/latent-flow.
@@
This paper is an extension of [song2023flow]. We build our model upon [song2023flow] in the high-level concept of using latent flow fields for modeling transformations, but we do have two substantial modifications: (1) our method uses spike-and-slab prior to avoid any sort of supervision; (2) we leverage Helmholtz decomposition for more expressive/flexible latent flows. These improvements could bring many concrete benefits. For example, the elimination of supervision greatly broadens the applicability of the method, thus making it applicable to real-world video understanding (e.g., CalMS [sun2021multi] and Cityscape [cordts2016cityscapes] in Sec. 5.5). Further, the slab component mimics the motion speed in natural videos and allows for the explicit control of transformation speeds, which is seldom studied in the literature of disentangled representation learning. Moreover, as discussed in Sec.5.4, Helmholtz decomposition allows each transformation to be associated with either a curl-free or divergence-free component, thereby offering improved interpretability of the learned transformation structure. Finally, to complement the empirical improvements, we also provide a formal identifiability argument grounded in sparse dictionary learning, as detailed in Sec. B of the supplementary material.
@@ 2.1 Disentangled/Equivariant Representation Learning
Disentanglement and equivariance are considered two important desiderata of representation learning [higgins2018towards]. The idea of learning disentangled representation was first studied by InfoGAN [chen2016infogan] and $\beta$-VAE [higgins2016beta]. InfoGAN [chen2016infogan] disentangles the latent space by maximizing the mutual information between a subset of latent dimensions and observations, while $\beta$-VAE [higgins2016beta] factorizes the posterior $q(\mathbf{z}|\mathbf{x})$ by penalizing the total correlation between the prior and variational posterior. Subsequent work following InfoGAN mainly focused on discovering different semantically interpretable directions in the latent space [goetschalckx2019ganalyze, jahanian2020steerability, voynov2020unsupervised, harkonen2020ganspace, zhu2020learning, peebles2020hessian, shen2021closed, wei2021orthogonal, zhu2021low, Tzelepis_2021_ICCV, zhu2022region, song2022orthogonal, oldfield2022panda, song2023latent, song2023householder]. Following $\beta$-VAE, many attempts have been made to encourage independence of the aggregated posterior through additional guidance [dilokthanakul2017deep, dupont2018learning, kumar2018variational, kim2018disentangling, chen2018isolating, jeong2019learning, yildiz2019ode2vae, ding2020guided, shao2020controlvae, locatello2020weakly, tai2022hyperbolic, estermann2023dava]. More recently, one line of research proposed to disentangle diffusion models by crafting compact low-dimensional latent spaces [kwon2022diffusion, park2023understanding, yang2023disdiff, wang2023infodiffusion]. Parallel developments in causal inference literature emphasize that disentanglement can be viewed as identifying independent causal mechanisms [spirtes2000causation, pearl2009causality, locatello2019challenging, yang2021causalvae]. Ke et al. [ke2019learning] demonstrated the potential for neural networks to learn causal models by interpreting unknown interventions. Kumar and Sinha [kumar2021disentangling] extended this idea by proposing methods for disentangling mixtures of unknown causal interventions, further bridging representation learning and causal inference. More recently, Song et al. [song2024causal] explored causal temporal representation learning, explicitly modeling nonstationary and sparse causal transitions in temporal data, thus advancing understanding of disentangled representations in dynamic scenarios.
Equivariant networks, on the other hand, are usually more strictly defined than disentanglement methods. Analytical approaches typically enforce neural network weights to explicitly respect the symmetry of group transformations [cohen2016group, cohen2017steerable, ravanbakhsh2017equivariance, worrall2017harmonic, worrall2019deep, finzi2020generalizing, hoogeboom2022equivariant]. However, as noted in the introduction, the transformation for which such analytic equivariance is possible is limited to certain groups like special orthogonal groups, and may not apply to real-world scenarios. To avoid this issue, a number of recent models have aimed to relax this constraint and instead learn approximately equivariant representations directly from data [diaconu2019learning, connor2021variational, klindt2021towards, keller2021topographic, song2023latent, song2023flow].
@@ 2.2 Sequential Disentanglement
Another closely related research branch of disentangled representations is sequential disentanglement [hsu_unsupervised2017, denton2017unsupervised, villegas2017decomposing, lisevae2018, tulyakov2018mocogan, zhu2020s3vae, bhagat2020disentangling, yamada2020disentangled, bai2021contrastively, han2021disentangled, tonekaboni2022decoupling, naiman2023operator, berman2023multi, nimrod2024sequential] where the disentangled representation learning techniques are applied to sequence data like video and audio. In the sequential case, latent variables are typically split into single static time-invariant codes that do not change over time and multiple dynamic time-varying components that describe the distinct motions in the sequence. Due to the static and dynamic assumptions, these methods have to use two sets of latent variables for modeling different components. Differently, we assume that the static identity information is given in the latent variable, and the dynamic sequential transformations are encoded in the latent flow fields. Further, these approaches achieve disentanglement of single latent dimensions through implicit KL regularization. By contrast, we leverage the sophisticated sparsity constraints to explicitly classify the transformations and factorize them into different latent flows.
@@ 2.3 Physical Inductive Biases in Deep Learning
The performance of deep learning models is heavily based on inductive biases. In recent years, an increasing amount of effort has developed to endow deep neural networks with physical priors and inductive biases (e.g., symmetries or conservation laws). Much attention has been focused on using neural networks to solve Partial Differentiable Equations (PDEs), such as Physics Informed Neural Networks (PINNs) [pinns] and other improved variants [hsieh2019learning, brandstetter2022message, richter2022neural, zeng2023competitive, bajaj2023recipes, akhound2023lie]. Other active research directions include handling input symmetries with aforementioned equivariant networks, building generative score-based denoising diffusion models using Fokker-Planck equations [ho2020denoising, song2020denoising, song2020score], and designing neural networks with Hamiltonian dynamics for improved generalization [greydanus2019hamiltonian, toth2019hamiltonian]. In this work, we leverage PINNs to place constraints on our latent flow fields such that they obey the assumptions of fluid-dynamic optimal transport and the Helmholtz decomposition, thereby increasing the expressivity while including valuable inductive biases.
@@ 3.2 Spike and Slab Priors
We model real-world video as a sparse combination of transformation primitives. To model this transition sparsity, we impose a spike and slab prior [mitchell1988bayesian] on the transformation variable ${\bm{g}}_{t}$ for generating the sequences. The distribution is factorized as follows:
@@
Fig. 3 displays the generated sequences of MNIST [lecun1998mnist] using spike priors. The variations align with natural videos – the transitions happen occasionally and smoothly.
@@ 3.3 Latent Prior Time Evolution
where ${\bm{v}}^{k}({\bm{z}})$ denotes the velocity field of the $k$’th latent flow. We use ${\bm{g}}_{t}$ to combine the vector fields linearly to model the possible multiple transformations. The conditional update can be derived from the change of variables formula [rezende2015variational, chen2019neural]:
@@ 4.1 Helmholtz Decomposed Latent Flows
By the Helmholtz decomposition [helmholtz1858integrale, helmholtz1867lxiii, abraham2012manifolds], a vector field $\mathbf{F}$ can be uniquely represented by the sum of two vector fields such that:
@@
Richter et al. [richter2022neural] proposed an approach to construct strict divergence-free vector fields. However, it requires computing the full Jacobian matrix at every step, which is memory-intensive and computationally slow. For faster computation, we use a PINN to approximate the vector field. Compared with prior work [song2023flow, song2023latent] which only includes the curl-free component $\mathbf{G}$, this parameterization allows for significantly increased flexibility in modeling periodic dynamics in the latent space. Furthermore, as will be illustrated later in Sec. 5.3, we expect that our model automatically learns to segregate periodic and non-periodic transformations into these two components.
@@ 4.2 Inference
Compared with the objective of a traditional VAE, our model additionally involves the time evolution of the priors and posteriors. As noted in Sec. 3.2, we set $p(\tilde{{\bm{g}}}_{t})$ to follow a Laplace distribution and impose multivariate Bernoulli distributions to $p({\bm{y}}_{1})$ and $p({\bm{y}}_{t}|{\bm{y}}_{t-1})$. The KL divergence on $q_{\gamma}({\bm{y}}_{t}|{\bm{x}}_{t},{\bm{x}}_{t-1})$ serves as regularization to encourage the sparsity of ${\bm{y}}_{t}$. That being said, the posterior $q_{\gamma}({\bm{y}}_{t}|{\bm{x}}_{t},{\bm{x}}_{t-1})$ learns to model the transformations using as few vector fields as possible, which naturally disentangles the input variations into distinct flow fields. We apply the Gumbel-Sigmoid trick [jang2017categorical] for the re-parameterization and sampling of $q_{\gamma}({{\bm{y}}}_{t}|{\bm{x}}_{t},{\bm{x}}_{t-1})$.
TABLE I: Equivariance error $\mathcal{E}_{k}$ and average log-likelihood $\log p({\bm{x}}_{t})$ on MNIST [lecun1998mnist].

### v3

@@ 1 Introduction
One early line of work in representation learning focused on ideas of redundancy reduction, believing that biological neural systems would naturally strive for an efficient code due to competitive pressures [6]. Building on this idea, the principles of sparsity and statistical independence of coding dimensions emerged as guidelines for learning such maximally efficient codes, eventually resulting in the frameworks of sparse coding [65] and independent component analysis [18]. Inspired by the fact that natural intelligence is embedded in a world where physical laws restrict observations to sequences of smooth transformations, these ideas of efficiency and sparsity were extended to include temporal dimensions. A seminal example is Slow Feature Analysis [96], a learning framework which assumes that individual latent variables are likely to change slowly over time. Models adhering to these principles were shown to learn invariances directly from data and uncover underlying generative factors if those factors had similar slow dynamics. Recent work has further shown that natural videos follow a specific sparse transition structure, meaning that the set of generative factors which describe a given input sequence is mostly constant over time with sparse transitions between which factors are active. Klindt et al. [52] then demonstrated that by building a model which incorporates this structure into its prior, it is possible to provably learn the true generative factors of video data in an unsupervised manner. While differing in implementation and methodology, all these frameworks appear in some sense to share the goals of learning meaningful, interpretable, ‘disentangled’, and controllable latent codes such that specific directions in the latent space corresponded to the independent factors which were responsible for generating the input data distribution.
@@
More recently, the concept of equivariance has emerged as a mathematical framework for learning highly structured and thereby ‘controllable’ latent representations [16]. Specifically, equivariant neural networks are built to explicitly respect the symmetries of the input domain in their output space. In such models, there are known predictable output transformations for given input transformations of interest. To date, these concepts form the foundation of some of the most precisely formalized definitions of ‘disentanglement’ in the literature [37, 15]. In prior work however, there is a relatively sharp divide between equivariant neural networks and models which are focused on disentanglement. Specifically, it is currently only known how to build networks which are equivariant with respect to known transformations which have a mathematical group structure. This includes traditional coordinate symmetries [17, 40], but is severely limiting when considering the types of natural image transformations that are typically explored in disentangled representation learning. One line of research has aimed to bridge this divide by learning ‘approximately equivariant’ models which are intended to learn these types of structured representations directly from data itself [49, 79]. However the vast majority of the models in this domain require at least some form of weak supervision of segmented sequences with only single transformations being observed.
@@
In this paper, we introduce a new modeling framework, denoted Sparse Transformation Analysis (STA), which takes inspiration from these foundational representation learning approaches, thereby yielding what we argue to be a uniquely structured yet flexible latent space which aligns with natural data statistics. Fig. 1 depicts the overview of our STA. The framework requires no supervision of input sequences, assuming only that the observed transformations from one timestep to the next match a sparse transition structure similar to that observed by [52]. Specifically, STA takes a generative modeling approach, asserting that generative factors should be represented by distributions over latent variables, and that these distributions should flow smoothly in the latent space in concert with the smooth flow of observations in the world. Furthermore, the framework posits that this flow should not be arbitrary, but can be represented as a sparse combination of learned flow field primitives. In alignment with notions of disentanglement and approximate equivariance, these flow field primitives can be seen as directions in latent space which correspond to observed input transformations. Unlike [52], STA allows for highly flexible latent dynamics for each transformation ‘direction’ by parameterizing each transformation’s flow field through the Helmholtz decomposition as a combination of curl-free and divergence-free components. The specific sparse combination of flow fields which are used to transform the latent distribution from one time-step to the next is treated as an unobserved latent variable with a multivariate history-dependent spike and slab prior [60], and inferred simultaneously with the other latent variables through amortized variational inference. In this way, this approach can be seen to combine ideas of sparse coding, slow feature analysis, and approximate equivariance, while still allowing deep neural network feature extractors to be leveraged in a relatively unconstrained manner.
@@
In the following, we will demonstrate that this framework yields the state of the art in unsupervised approximate equivariance, as quantified through a measured equivariance error, and further that our method yields the highest likelihood on the test set in the unsupervised setting. As desired, we additionally observe that the model automatically learns to separate observed transformations into independent flow fields, and that these latent flows can be flexibly combined or switched during traversal. By tuning the magnitude of these flow fields, our model also has precise control of the transformation speed. Further, when slightly modifying the spike component to two separate controls, our method can learn to segregate latent symmetries and invariances into the two distinct components of the vector fields. The decomposition of latent flow fields is coherent with the categories of input transformations. Besides simple experiments on toy datasets, we further validate our STA on use cases of real-world video analysis, including movements of robot arms [62], lighting changes of indoor scenes [62], behavior videos of social agents [85], and ego-centric autonomous driving videos [20]. Our method can identify a wide range of independent motions in the video sequences. Ultimately, we present this model as a natural next step in the development of unsupervised approximately equivariant representation learning algorithms. Code is publicly available at https://github.com/KingJamesSong/latent-flow.
@@
This paper is an extension of [79]. We build our model upon [79] in the high-level concept of using latent flow fields for modeling transformations, but we do have two substantial modifications: (1) our method uses spike-and-slab prior to avoid any sort of supervision; (2) we leverage Helmholtz decomposition for more expressive/flexible latent flows. These improvements could bring many concrete benefits. For example, the elimination of supervision greatly broadens the applicability of the method, thus making it applicable to real-world video understanding (e.g., CalMS [85] and Cityscape [20] in Sec. 5.5). Further, the slab component mimics the motion speed in natural videos and allows for the explicit control of transformation speeds, which is seldom studied in the literature of disentangled representation learning. Moreover, as discussed in Sec.5.4, Helmholtz decomposition allows each transformation to be associated with either a curl-free or divergence-free component, thereby offering improved interpretability of the learned transformation structure. Finally, to complement the empirical improvements, we also provide a formal identifiability argument grounded in sparse dictionary learning, as detailed in Sec. B of the supplementary material.
@@ 2.1 Disentangled/Equivariant Representation Learning
Disentanglement and equivariance are considered two important desiderata of representation learning [37]. The idea of learning disentangled representation was first studied by InfoGAN [14] and $\beta$-VAE [38]. InfoGAN [14] disentangles the latent space by maximizing the mutual information between a subset of latent dimensions and observations, while $\beta$-VAE [38] factorizes the posterior $q(\mathbf{z}|\mathbf{x})$ by penalizing the total correlation between the prior and variational posterior. Subsequent work following InfoGAN mainly focused on discovering different semantically interpretable directions in the latent space [31, 45, 93, 34, 106, 68, 75, 95, 104, 90, 105, 81, 64, 80, 82]. Following $\beta$-VAE, many attempts have been made to encourage independence of the aggregated posterior through additional guidance [23, 26, 54, 50, 13, 47, 102, 24, 74, 59, 86, 29]. More recently, one line of research proposed to disentangle diffusion models by crafting compact low-dimensional latent spaces [55, 66, 101, 94]. Parallel developments in causal inference literature emphasize that disentanglement can be viewed as identifying independent causal mechanisms [84, 67, 58, 100]. Ke et al. [48] demonstrated the potential for neural networks to learn causal models by interpreting unknown interventions. Kumar and Sinha [53] extended this idea by proposing methods for disentangling mixtures of unknown causal interventions, further bridging representation learning and causal inference. More recently, Song et al. [77] explored causal temporal representation learning, explicitly modeling nonstationary and sparse causal transitions in temporal data, thus advancing understanding of disentangled representations in dynamic scenarios.
Equivariant networks, on the other hand, are usually more strictly defined than disentanglement methods. Analytical approaches typically enforce neural network weights to explicitly respect the symmetry of group transformations [16, 17, 70, 97, 98, 30, 40]. However, as noted in the introduction, the transformation for which such analytic equivariance is possible is limited to certain groups like special orthogonal groups, and may not apply to real-world scenarios. To avoid this issue, a number of recent models have aimed to relax this constraint and instead learn approximately equivariant representations directly from data [22, 19, 52, 49, 80, 79].
@@ 2.2 Sequential Disentanglement
Another closely related research branch of disentangled representations is sequential disentanglement [42, 21, 92, 57, 89, 107, 9, 99, 4, 33, 87, 61, 8, 63] where the disentangled representation learning techniques are applied to sequence data like video and audio. In the sequential case, latent variables are typically split into single static time-invariant codes that do not change over time and multiple dynamic time-varying components that describe the distinct motions in the sequence. Due to the static and dynamic assumptions, these methods have to use two sets of latent variables for modeling different components. Differently, we assume that the static identity information is given in the latent variable, and the dynamic sequential transformations are encoded in the latent flow fields. Further, these approaches achieve disentanglement of single latent dimensions through implicit KL regularization. By contrast, we leverage the sophisticated sparsity constraints to explicitly classify the transformations and factorize them into different latent flows.
@@ 2.3 Physical Inductive Biases in Deep Learning
The performance of deep learning models is heavily based on inductive biases. In recent years, an increasing amount of effort has developed to endow deep neural networks with physical priors and inductive biases (e.g., symmetries or conservation laws). Much attention has been focused on using neural networks to solve Partial Differentiable Equations (PDEs), such as Physics Informed Neural Networks (PINNs) [69] and other improved variants [41, 10, 72, 103, 5, 3]. Other active research directions include handling input symmetries with aforementioned equivariant networks, building generative score-based denoising diffusion models using Fokker-Planck equations [39, 76, 78], and designing neural networks with Hamiltonian dynamics for improved generalization [32, 88]. In this work, we leverage PINNs to place constraints on our latent flow fields such that they obey the assumptions of fluid-dynamic optimal transport and the Helmholtz decomposition, thereby increasing the expressivity while including valuable inductive biases.
@@ 3.2 Spike and Slab Priors
We model real-world video as a sparse combination of transformation primitives. To model this transition sparsity, we impose a spike and slab prior [60] on the transformation variable ${\bm{g}}_{t}$ for generating the sequences. The distribution is factorized as follows:
@@
Fig. 3 displays the generated sequences of MNIST [56] using spike priors. The variations align with natural videos – the transitions happen occasionally and smoothly.
@@ 3.3 Latent Prior Time Evolution
where ${\bm{v}}^{k}({\bm{z}})$ denotes the velocity field of the $k$’th latent flow. We use ${\bm{g}}_{t}$ to combine the vector fields linearly to model the possible multiple transformations. The conditional update can be derived from the change of variables formula [71, 12]:
@@ 4.1 Helmholtz Decomposed Latent Flows
By the Helmholtz decomposition [35, 36, 1], a vector field $\mathbf{F}$ can be uniquely represented by the sum of two vector fields such that:
@@
Richter et al. [72] proposed an approach to construct strict divergence-free vector fields. However, it requires computing the full Jacobian matrix at every step, which is memory-intensive and computationally slow. For faster computation, we use a PINN to approximate the vector field. Compared with prior work [79, 80] which only includes the curl-free component $\mathbf{G}$, this parameterization allows for significantly increased flexibility in modeling periodic dynamics in the latent space. Furthermore, as will be illustrated later in Sec. 5.3, we expect that our model automatically learns to segregate periodic and non-periodic transformations into these two components.
@@ 4.2 Inference
Compared with the objective of a traditional VAE, our model additionally involves the time evolution of the priors and posteriors. As noted in Sec. 3.2, we set $p(\tilde{{\bm{g}}}_{t})$ to follow a Laplace distribution and impose multivariate Bernoulli distributions to $p({\bm{y}}_{1})$ and $p({\bm{y}}_{t}|{\bm{y}}_{t-1})$. The KL divergence on $q_{\gamma}({\bm{y}}_{t}|{\bm{x}}_{t},{\bm{x}}_{t-1})$ serves as regularization to encourage the sparsity of ${\bm{y}}_{t}$. That being said, the posterior $q_{\gamma}({\bm{y}}_{t}|{\bm{x}}_{t},{\bm{x}}_{t-1})$ learns to model the transformations using as few vector fields as possible, which naturally disentangles the input variations into distinct flow fields. We apply the Gumbel-Sigmoid trick [46] for the re-parameterization and sampling of $q_{\gamma}({{\bm{y}}}_{t}|{\bm{x}}_{t},{\bm{x}}_{t-1})$.
TABLE I: Equivariance error $\mathcal{E}_{k}$ and average log-likelihood $\log p({\bm{x}}_{t})$ on MNIST [56].